# Teaching LLMs to See and Guide: Context-Aware Real-Time Assistance in Augmented Reality

### v1

@@ Abstract
The growing adoption of augmented and virtual reality (AR/VR) technologies in industrial training and on-the-job assistance has created new opportunities for intelligent, context-aware support systems. As workers perform complex tasks guided by AR/VR, these devices capture rich streams of multimodal data—including gaze, hand actions, and task progression—that can reveal user intent and task state in real time. Leveraging this information effectively remains a major challenge. In this work, we present a context-aware large language model (LLM) assistant that integrates diverse data modalities—such as hand actions, task steps, and dialogue history—into a unified framework for real-time question answering. To systematically study how context influences performance, we introduce an incremental prompting framework, where each model version receives progressively richer contextual inputs. Using the HoloAssist dataset, which records AR-guided task executions, we evaluate how each modality contributes to the assistant’s effectiveness. Our experiments show that incorporating multimodal context significantly improves the accuracy and relevance of responses. These findings highlight the potential of LLM-driven multimodal integration to enable adaptive, intuitive assistance for AR/VR-based industrial training and assistance.
@@ I Introduction
Augmented Reality (AR) is rapidly gaining traction in industrial training and on-the-job assistance, where workers perform complex, safety-critical tasks that demand precision and adaptability [1, 2, 3, 4]. By overlaying digital information directly within the user’s field of view, AR enhances productivity, reduces cognitive load, accelerates skill acquisition, and minimizes operational errors across domains ranging from industrial maintenance to surgical procedures. As AR and virtual reality (VR) technologies become increasingly embedded in workplaces, they generate rich streams of multimodal data—including gaze, hand motions, dialogue, and task progression—that can reveal valuable insights into user intent and task context in real time. Leveraging this information effectively can transform AR from a passive display tool into an intelligent, adaptive assistant capable of understanding what the user is doing and anticipating what they need next [5, 6, 7, 8]. At the same time, recent advances in large language models (LLMs) such as GPT have revolutionized machine understanding and interaction. These models excel at reasoning across modalities, adapting to context, and engaging in natural, task-specific dialogue [9, 10, 11, 12, 13]. Their success in question answering, content summarization, and instructional guidance highlights their potential to support real-time decision-making and knowledge transfer in AR environments.
@@
  • We propose a context-aware framework for AR-based task guidance that integrates LLMs with multimodal AR sensor inputs—such as hand actions, egocentric video, and speech—to deliver adaptive, human-like, and real-time assistance. This integration enables contextually-aware guidance that dynamically adapts to the user’s ongoing actions and environment.
@@ III-A1 Model 1—Minimal Context (Baseline)
This model receives only high-level task metadata (e.g., task type, duration, or goal). The input excludes any procedural or temporal context, simulating scenarios where the assistant has limited background information. Implementation details:
@@ III-A2 Model 2—Task Overview Context
This configuration adds structured task-step information, representing the full sequence of operations required to complete the task. This gives the model a procedural overview, enabling step-aware reasoning. Additional inputs: structured step list (e.g., Step 1N1–N with brief descriptions). Implementation details:
@@ III-A3 Model 3—Real-Time Step Context
Building upon Model 2, this version integrates temporal awareness by identifying the user’s current task step at the time a query is made (time $T$). This allows the model to tailor its response specifically to the user’s in-progress activity. Additional inputs: current step index or label, task progress indicator, or timestamp alignment. Implementation details:
@@ III-A4 Model 4—Full Multimodal Context
The final configuration extends the context further by incorporating hand action data, encoded as verb–noun pairs (e.g., “pick tool,” “measure part”) extracted from real-time motion segmentation. These actions, typically spanning 1–2 seconds, provide fine-grained behavioral cues that enrich situational understanding. Additional inputs: hand action segments, action-confidence scores, and optional dialogue snippets. Implementation details:
@@ V-B LLM-as-a-judge
To complement the quantitative metrics, we employed a LLLM as an automatic judge to assess response quality more holistically. The judge model was provided with all contextual information available up to the point each question was asked, including task description, current task step, hand actions data, and conversation history, and was instructed to evaluate the four model-generated responses based on correctness, completeness, contextual relevance, and clarity. This LLM-based evaluation enables the capture of semantic and pragmatic aspects that lexical similarity metrics cannot reliably measure. To mitigate potential positional bias in the judging process (i.e., the tendency of an LLM to favor responses appearing earlier in the list), we randomized the order of the four model outputs across four different permutations: 1234, 3421, 4321, and 2143. While exhaustively evaluating all $4!$ orderings would be computationally expensive, we selected four representative permutations that allow each model to appear in every possible position, which we believe sufficiently mitigates positional bias. The final score for each model represents the average across these permutations. Each model received individual 0–5 ratings on all dimensions, along with an overall comparison summary identifying the most contextually appropriate response. The full judging prompt template is provided in Appendix B-A.
@@ VI Discussion
In this study, we developed a comprehensive framework for context-aware, real-time LLM assistants designed to support task guidance in AR environments. The observed improvement in response quality with the addition of contextual information suggests that fine-grained physical and semantic context significantly enhances LLM assistants performance. While our analysis was conducted using annotated tasks from the HoloAssist dataset, which may not fully reflect the variability of real-world scenarios, our findings lay a strong foundation for future deployment.
As the next step, we aim to extend this framework to real-world AR and VR task environments. We plan to collect a dataset of human-human collaborative interactions and establish a benchmark for context-aware LLM assistants in this domain. In collaboration with other researchers in our lab working on state-of-the-art hand action recognition and step detection models, we will apply these techniques to live scenarios to extract the essential contextual inputs required by the assistant in real time, Figure 14.
We also intend to enrich the contextual input space by incorporating additional signals such as gaze data, video frames, and video captions. While the current work focuses primarily on question answering, we aim to expand the assistant’s capabilities to provide proactive task guidance when users are stuck without initiating a query, to trigger specific robot behaviors in human–robot interaction tasks based on context, and to support VR environments in addition to AR.
Overall, this work represents an important step toward building intelligent, responsive, and contextually grounded mixed-reality systems enabled by LLMs.
@@ Appendix B Models’ Prompt Templates
B-A LLM-as-a-judge Prompt Template

### v2

@@ Abstract
The growing adoption of augmented and virtual reality (AR/VR) technologies in industrial training and on-the-job assistance has created new opportunities for intelligent, context-aware support systems. As workers perform complex tasks guided by AR/VR, these devices capture rich streams of multimodal data—including gaze, hand actions, and task progression—that can reveal user intent and task state in real time. Leveraging this information effectively remains a major challenge. In this work, we present a context-aware large language model (LLM) assistant that integrates diverse data modalities, such as hand actions, task steps, and dialogue history, into a unified framework for real-time question answering. To systematically study how context influences performance, we introduce an incremental prompting framework, where each model version receives progressively richer contextual inputs. Using the HoloAssist dataset, which records AR-guided task executions, we evaluate how each modality contributes to the assistant’s effectiveness. Our experiments show that incorporating multimodal context significantly improves the accuracy and relevance of responses. These findings highlight the potential of LLM-driven multimodal integration to enable adaptive, intuitive assistance for AR/VR-based industrial training and assistance.
@@ I Introduction
Augmented Reality (AR) is rapidly gaining traction in industrial training and on-the-job assistance, where workers perform complex, safety-critical tasks that demand precision and adaptability [1, 2, 3, 4]. By overlaying digital information directly within the user’s field of view, AR enhances productivity, reduces cognitive load, accelerates skill acquisition, and minimizes operational errors across domains ranging from industrial maintenance to surgical procedures. As AR and virtual reality (VR) technologies become increasingly embedded in workplaces, they generate rich streams of multimodal data, including gaze, hand motions, dialogue, and task progression, that can reveal valuable insights into user intent and task context in real time. Leveraging this information effectively can transform AR from a passive display tool into an intelligent, adaptive assistant capable of understanding what the user is doing and anticipating what they need next [5, 6, 7, 8]. At the same time, recent advances in large language models (LLMs) such as GPT have revolutionized machine understanding and interaction. These models excel at reasoning across modalities, adapting to context, and engaging in natural, task-specific dialogue [9, 10, 11, 12, 13]. Their success in question answering, content summarization, and instructional guidance highlights their potential to support real-time decision-making and knowledge transfer in AR environments.
@@
  • We propose a context-aware framework for AR-based task guidance that integrates LLMs with multimodal AR sensor inputs, such as hand actions, egocentric video, and speech, to deliver adaptive, human-like, and real-time assistance. This integration enables contextually-aware guidance that dynamically adapts to the user’s ongoing actions and environment.
@@ III-A1 Model 1—Minimal Context (Baseline)
This model receives only high-level task metadata (e.g., task type, duration, or goal). The input excludes any procedural or temporal context, simulating scenarios where the assistant has limited background information.
@@ III-A2 Model 2—Task Overview Context
This configuration adds structured task-step information, representing the full sequence of operations required to complete the task. This gives the model a procedural overview, enabling step-aware reasoning. Additional inputs: structured step list (e.g., Step 1N1–N with brief descriptions).
@@ III-A3 Model 3—Real-Time Step Context
Building upon Model 2, this version integrates temporal awareness by identifying the user’s current task step at the time a query is made (time $T$). This allows the model to tailor its response specifically to the user’s in-progress activity. Additional inputs: current step index or label, task progress indicator, or timestamp alignment.
@@ III-A4 Model 4—Full Multimodal Context
The final configuration extends the context further by incorporating hand action data, encoded as verb–noun pairs (e.g., “pick tool,” “measure part”) extracted from real-time motion segmentation. These actions, typically spanning 1–2 seconds, provide fine-grained behavioral cues that enrich situational understanding. Additional inputs: hand action segments, action-confidence scores, and optional dialogue snippets.
@@ V Results
This section compares the four models using lexical metrics, LLM-as-a-judge scoring, and human evaluations, and further analyzes LLM–human agreement and the framework’s sensitivity to individual elements through ablation studies.
@@ V-B LLM-as-a-Judge
To complement the quantitative metrics, we employed a LLM as an automatic judge to assess response quality more holistically. The judge model was provided with all contextual information available up to the point each question was asked, including task description, current task step, hand actions data, and conversation history, and was instructed to evaluate the four model-generated responses based on correctness, completeness, contextual relevance, and clarity. This LLM-based evaluation enables the capture of semantic and pragmatic aspects that lexical similarity metrics cannot reliably measure. To mitigate potential positional bias in the judging process (i.e., the tendency of an LLM to favor responses appearing earlier in the list), we randomized the order of the four model outputs across four different permutations: 1234, 3421, 4321, and 2143. While exhaustively evaluating all $4!$ orderings would be computationally expensive, we selected four representative permutations that allow each model to appear in every possible position, which we believe sufficiently mitigates positional bias. The final score for each model represents the average across these permutations. Each model received individual 0–5 ratings on all dimensions, along with an overall comparison summary identifying the most contextually appropriate response. The full judging prompt template is provided in Appendix C.
@@ VI Discussion
In this study, we introduced a comprehensive framework for context-aware, real-time LLM assistants designed to support task guidance in AR environments. The observed improvements in response quality with the addition of contextual information indicate that fine-grained physical and semantic context substantially enhances the performance of LLM-based assistants. While our analysis was conducted using annotated tasks from the HoloAssist dataset—which may not fully capture the variability of real-world scenarios—our findings provide a strong foundation for future deployment and exploration. Yet, several fundamental research directions remain:
  • Deployment in real-world AR/VR environments: Future research should extend this framework to real-world AR and VR task environments. To this end, we plan to collect and annotate a new dataset of human–human collaborative interactions and establish a benchmark for context-aware LLM assistants in human–robot collaboration settings. These will involve more cognitively challenging and physically demanding real-world tasks, such as assembly and maintenance. In collaboration with researchers developing state-of-the-art hand action recognition and step detection models, we will integrate these techniques into live scenarios to extract essential contextual cues required by the assistant in real time (see Figure 14).
  • Expansion of modalities and capabilities: Future studies may also enrich the contextual input space by incorporating additional behavioral signals such as gaze trajectories, ego-/exocentric video streams, and video captions. While the current work focuses primarily on question answering, we aim to expand the LLM assistant’s capabilities toward proactive task guidance—detecting when a user is stuck and offering help without explicit queries—and toward initiating specific robot behaviors in human–robot interaction tasks. Additionally, we plan to extend the framework to support fully immersive VR environments alongside AR.
  • Foundation models for behavioral understanding: A fundamentally challenging and potentially transformative direction is to develop multimodal foundation models (LLM/VLM/VLA) capable of directly interpreting raw behavioral cues (e.g., gaze, gestures, body pose) and physiological signals (e.g., EEG, heart rate variability) to adaptively guide users and proactively intervene during complex task execution. Our current implementation—even when deployed in AR for real-time question answering—relies on intermediate, interpretable inferences for activity understanding (e.g., step recognition, error detection) and user state estimation (e.g., cognitive load, visual attention). Future research should explore end-to-end inference models capable of engaging in mixed-initiative dialogue grounded directly in raw, multimodal data from AR/VR headsets and wearable sensors.
  • Privacy, safety, and overreliance on technology: The proposed framework depends on continuous capture and monitoring of behavioral data, which may compromise user privacy and therefore limit adoption. New approaches are needed to ensure secure, anonymized data collection, processing, and disposal. Moreover, LLMs can make errors—sometimes with serious implications. Thus, it is essential to design safeguards that prevent harmful outcomes and to implement dynamic mechanisms for shared control between human and AI. Additionally, continuous exposure to assistive technologies may lead to overdependence—just like how GPS navigation can limit our ability to navigate independently. To prevent these undesirable consequences, future systems should adaptively scaffold and fade guidance based on user performance, expertise, and interaction history to preserve and strengthen human skills.
Overall, this work represents an important step toward building intelligent, responsive, and contextually grounded AR/VR systems enabled by large language models.
@@ Appendix B Models’ Prompt Templates
Appendix C LLM-as-a-judge Prompt Template